OpenStudy (anonymous):

how many solutions does 6t-3=3t+6 have

OpenStudy (midhun.madhu1987):

6t - 3 = 3t + 6 6t - 3t = 6 + 3 Can you solve it ??

OpenStudy (anonymous):

If 3t = 9, then one t = 3 .

OpenStudy (dmndlife24):